# 11footnotetext: Corresponding author

### v1

@@ 4.3 Ablation Study and Analyses
Visionary-R1 64.6 22.7 62.9 87.6
Figure 5: Visualization of learning curves for different KL coefficients (top) and output examples (bottom).
Table 3: Results of using different KL coefficients. Dynamic strategies (i.e., linear decay and cosine annealing) achieve significantly better results in Visionary-R1, with cosine annealing being the optimal choice.
Method Strategy MathVista MathVision
Visionary-R1 Static (0.04) 60.9 19.3
Static (0.008) 60.7 18.7
Linear 63.4 22.4
Cosine 64.6 22.7
GRPO Static (0.04) 59.0 18.2
Cosine 59.6 18.4
@@
We conduct an ablation study to evaluate the effectiveness of each component in Visionary-R1. Specifically, we start from the GRPO model and incrementally add the caption output format and the caption reward $r_{c}$. Instead of using the compiled 272.6K training data, we use individual datasets to save computation. Specifically, we perform two sets of experiments on different types of datasets (to ensure diversity): 1) training on ChartQA and testing on MathVista and MathVision, and 2) training on A-OKVQA and testing on MMStar and MMBench. Tab. 2 shows the results, which clearly demonstrate the effectiveness of the caption output and the caption reward. Fig. 4 further illustrates the differences in the outputs of different models.
@@
To better understand why the KL coefficient makes such a huge impact, we dig into several key metrics logged during training, i.e., output length, the format reward, and the caption reward. The full training processes are shown in Fig. 5 (top). When setting the KL coefficient to 0.04, which has been widely adopted as a standard practice in the literature, the output length rapidly climbs up and reaches an unreasonably high value at around 700 steps, and then falls back to the normal level at 100 tokens; meanwhile, both the format reward and caption reward decline drastically as the output length shoots up to an abnormal value, meaning that the model has collapsed in the middle of training. The model collapse is more clear in Fig. 5 (bottom): the model generates long but completely meaningless reasoning tokens. When using a smaller value of 0.008, we encounter the reward hacking issue [37]: the model mistakenly generates a short reasoning chain at the caption place (which is supposed to contain a description about the image) while producing zero token in between <think></think>. This suggests that the model cheats in order to gain a higher accuracy reward and as a result the reasoning capabilities are not generalizable. The use of either linear decay or cosine annealing can effectively alleviate this issue.
5 Conclusion, Limitations, and Future Work
This paper reveals the shortcut learning problem encountered when applying RL to VLMs. Unlike LLMs, VLMs are more difficult to train for reasoning without using annotated data. Visionary-R1, despite using CoT-free question-answer pairs, demonstrates strong performance on challenging visual reasoning benchmarks, surpassing strong commercial AI models that mostly likely benefit from larger-scale, higher-quality training data. The results indicate that understanding image context through captioning is essential for enhancing reasoning for VLMs. Moreover, the results also highlight the importance of the KL coefficient, which should be dynamically tuned to stabilize RL training. We believe the finding of the cosine annealing strategy could be applied more broadly to other RL applications. In terms of limitations, the experiments are only based on a 3B model due to tight budget on compute. We believe that the effectiveness of RL training can be significantly amplified by using larger models. Investigation on larger-scale models is left as future work.
@@ A.1 Complete list of training data
Tab. 4 shows the complete training data, which aggregates 11 popular question-answer datasets and covers a wide range of visual formats and tasks, e.g., A-OKVQA [38] and TextVQA [39] for general scene understanding, ChartQA [40] and RoBUT SQA [41] for chart understanding, GeoQA+ [42] for mathematical problem-solving, and DocVQA [43] for document processing.
@@
A-OKVQA [38] 17.1K Multi-choice General Scene
ChartQA [40] 28.3K Open-text+Num Chart
AI2D [44] 15.5K Multi-choice Diagram
ScienceQA [45] 6.2K Multi-choice Scene + Chart
GeoQA+ [42] 12.1K Multi-choice Math
DocVQA [43] 39.5K Open-text Document
CLEVR-Math [46] 32.6K Num 3D
Icon-QA [47] 29.9K Multi-choice Diagram
TabMWP [48] 23.1K Open-text+Num Table
RoBUT SQA [41] 34.1K Open-text+Num Chart
TextVQA [39] 34.6K Multi-choice General Scene
@@ A.3 Caption Reward Prompt
A.4 Visualization of the Visionary-R1 Output

### v3

@@ 4.3 Ablation Study and Analyses
GRPO+Caption+Length Reward 62.0 20.3 59.6 85.2
GRPO+Caption+Caption Reward 64.6 22.7 62.9 87.6
Figure 5: Visualization of curves for different KL coefficients (top) and output examples (bottom).
Table 3: Results of using different KL coefficients. Dynamic strategies (i.e., linear decay and cosine annealing) achieve significantly better results, with cosine annealing being the optimal choice.
Method Strategy MathVista MathVision MMStar MMBench
Visionary-R1 Static (0.04) 60.9 19.3 54.2 82.6
Static (0.008) 60.7 18.7 56.0 82.7
Linear 63.4 22.4 60.4 84.6
Cosine 64.6 22.7 61.6 85.5
GRPO Static (0.04) 59.0 18.2 48.1 80.4
Cosine 59.6 18.4 46.6 80.9
@@
We conduct an ablation study to evaluate the effectiveness of each component in Visionary-R1. Specifically, we start from the GRPO model and incrementally add the caption output format and the caption reward $r_{c}$. Instead of using the compiled 272.6K training data, we use individual datasets to save computation. Specifically, we perform two sets of experiments on different types of datasets (to ensure diversity): 1) training on ChartQA and testing on MathVista and MathVision, and 2) training on A-OKVQA and testing on MMStar and MMBench. Tab. 2 shows the results, which clearly demonstrate the effectiveness of the caption output and the caption reward. Fig. 4 further illustrates the differences in the outputs of different models. We also experiment with a simple length reward to encourage more detailed captions, but find that this superficial approach merely increases redundancy and reduces model performance. This further highlights the effectiveness of our caption reward strategy.
@@
To better understand why the KL coefficient makes such a huge impact, we dig into several key metrics logged during training, i.e., output length, the format reward, and the caption reward. The full training processes are shown in Fig. 5 (top). When setting the KL coefficient to 0.04, which has been widely adopted as a standard practice in the literature, the output length rapidly climbs up and reaches an unreasonably high value at around 700 steps, and then falls back to the normal level at 100 tokens; meanwhile, both the format reward and caption reward decline drastically as the output length shoots up to an abnormal value, meaning that the model has collapsed in the middle of training. The model collapse is more clear in Fig. 5 (bottom): the model generates long but completely meaningless reasoning tokens. When using a smaller value of 0.008, we encounter the reward hacking issue (Stiennon et al., 2020): the model mistakenly generates a short reasoning chain at the caption place (which is supposed to contain a description about the image) while producing zero token in between <think></think>. This suggests that the model cheats in order to gain a higher accuracy reward and as a result the reasoning capabilities are not generalizable. The use of either linear decay or cosine annealing can effectively alleviate this issue.
5 Conclusion and Future Work
This paper reveals the shortcut learning problem encountered when applying RL to VLMs. Unlike LLMs, VLMs are more difficult to train for reasoning without using annotated data. Visionary-R1, despite using CoT-free question-answer pairs, demonstrates strong performance on challenging visual reasoning benchmarks, surpassing strong commercial AI models that mostly likely benefit from larger-scale, higher-quality training data. The results indicate that understanding image context through captioning is essential for enhancing reasoning for VLMs. Moreover, the results also highlight the importance of the KL coefficient, which should be dynamically tuned to stabilize RL training. We believe the finding of the cosine annealing strategy could be applied more broadly to other RL applications. We believe that the effectiveness of RL training can be significantly amplified by using larger models. Investigation on larger-scale models is left as future work.
@@ A.1 Complete list of training data
Tab. 4 shows the complete training data, which aggregates 11 popular question-answer datasets and covers a wide range of visual formats and tasks, e.g., A-OKVQA (Schwenk et al., 2022) and TextVQA (Singh et al., 2019) for general scene understanding, ChartQA (Masry et al., 2022) and RoBUT SQA (Zhao et al., 2023) for chart understanding, GeoQA+ (Cao & Xiao, 2022) for mathematical problem-solving, and DocVQA (Mathew et al., 2021) for document processing.
@@
A-OKVQA (Schwenk et al., 2022) 17.1K Multi-choice General Scene
ChartQA (Masry et al., 2022) 28.3K Open-text+Num Chart
AI2D (Kembhavi et al., 2016) 15.5K Multi-choice Diagram
ScienceQA (Lu et al., 2022a) 6.2K Multi-choice Scene + Chart
GeoQA+ (Cao & Xiao, 2022) 12.1K Multi-choice Math
DocVQA (Mathew et al., 2021) 39.5K Open-text Document
CLEVR-Math (Lindström & Abraham, 2022) 32.6K Num 3D
Icon-QA (Lu et al., 2021) 29.9K Multi-choice Diagram
TabMWP (Lu et al., 2022b) 23.1K Open-text+Num Table
RoBUT SQA (Zhao et al., 2023) 34.1K Open-text+Num Chart
TextVQA (Singh et al., 2019) 34.6K Multi-choice General Scene
@@ A.3 Caption Reward Prompt
A.4 Additional Experimental Results
Benchmark Evaluation
Table 5 presents results on three challenging and diverse benchmarks: MMMUPro (Yue et al., 2024), MMStar (Chen et al., 2024a), and CV-Bench (Tong et al., 2024). Unlike prior methods, which often display inconsistent performance and even significant regressions on certain datasets, our approach, Visionary-R1, achieves consistent improvements across all benchmarks. This stability indicates that the model’s learned reasoning ability extends beyond dataset-specific adaptations, reflecting a more general and dependable form of multimodal reasoning.
Table 5: Comparison results on the additional three challenging visual benchmarks. Visionary-R1 achieves stable improvements across all datasets.
Methods Size MMMUPro MMStar CV‑Bench‑2D CV‑Bench‑3D
Base Model 7B 42.5 48.0 69.8 54.2
R1‑VL 7B 29.1 (-13.4) 60.0 (+12.0) 67.2 (-2.6) 65.9(11.7)
Base Model 7B 38.3 63.9 74.1 72.6
R1‑Onevision 7B 21.9 (-16.4) 59.1 (-4.8) 34.2 (-39.9) 20.1 (-52.5)
Base Model 3B 31.6 52.4 72.6 71.1
Visionary‑R1 3B 34.0 (+2.4) 61.5 (+9.1) 74.4 (+1.8) 74.0 (+2.9)
Model Scale Up Result
To evaluate the scalability of our approach in model scale, we conducted experiments using Qwen2.5-VL-7B as the base model and the A-OKVQA dataset (17.1K samples) for training. As shown in the Tab. 6, our method consistently outperforms the base model across all benchmarks for both the 3B and 7B model variants. These results provide strong evidence for the effectiveness and generalizability of our method.
Table 6: Experimental results of model scaling. The 17K data corresponds to training with the A-OKVQA dataset.
Methods Size RL Data MathVista MathVision MMStar MMBench
Base Model 3B - 61.5 19.1 52.4 82.1
Visionary-R1 3B 17K 62.5 20.5 62.9 87.6
Base Model 7B - 68.1 22.5 63.2 83.9
Visionary-R1 7B 17K 70.2 24.4 66.7 89.5
Hyper-parameter Analysis
For the GRPO method, we simply followed the original recipe by setting equal weights for the accuracy and format rewards. For the caption reward, we tried 0.1 and 0.5 as the balancing weight. The results are shown in the Tab. 7. The conclusion is that there is no significant difference between these two sets of results. To maintain training stability, we chose 0.1 and recommend that practitioners also use this setting.
Table 7: Results of hyper-parameter analysis for the reward weight $\alpha$.
Weight $\alpha$ MathVista MathVision MMStar MMBench
0.1 63.5 20.7 60.4 82.1
0.5 63.1 20.8 60.4 82.5
A.5 The Use of Large Language Models (LLMs)
No LLMs were used in any part of the research ideation, writing, or editing processes for this paper. All content was independently completed by the authors.
A.6 Visualization of the Visionary-R1 Output